# Syn3DWound: A Synthetic Dataset for 3D Wound Bed Analysis

### v1

@@ 3.1 Baseline systems and evaluation metrics
3D wound reconstruction: A 3D reconstruction algorithm estimates the 3D geometry of an object from a collection of 2D images. The prevailing methods in the literature rely on standard projective geometry techniques such as structure-from-motion and multiview stereopsis  [23, 12, 13]. However, new deep learning approaches for 3D scene rendering (e.g. Neural Radiance Fields (NeRF) [24]), are becoming very competitive. In this paper, we conduct a comparative analysis of two prominent open-source tools for 3D reconstruction: COLMAP [25] and Meshroom [26]. We also assess the performance of NeusFacto, a NeRF model tailored for surface extraction from the open-source SDFStudio toolbox [24].
We compared the 3D reconstructed meshes with the ground-truth synthetic mesh, after alignment using three steps: i) align the camera positions of the ground-truth data with those estimated by the frameworks (by solving a Procrustes problem [27]); ii) crop both meshes using the ground-truth 3D mask for wound bed segmentation, followed by fine alignment using the Iterative Closest Point (ICP) algorithm (applied only to the cropped meshes); ii) apply the transformations to the original meshes, followed by cropping the wound area again to report performance on the wound area only. In Table 2, we report the Average Symmetric Distance (ASD), Hausdorff Distance (HD90), and Normal Consistency (NC) metrics.
@@
3D wound bed segmentation: We introduce a 3D wound segmentation technique that assigns 2-dimensional labels to different regions of the reconstructed 3D models. We used a Meshroom-based texturing algorithm [29] to project a set of 2D wound segmentation masks onto 3D mesh vertices labeled as background and wound bed.
@@ 3.2 Results and discussion
3D wound reconstruction: Quantitative results for 3D wound reconstruction are reported in Table 2. In our experiment, COLMAP demonstrates superior surface accuracy, while the performance of the Neural rendering-based method is nearly comparable.
@@
3D wound segmentation: Table 4 compares 3D wound segmentation results with ground truth using previously described metrics. Notably, for the second sample, incorporating a higher number of 2D segmentation maps enhances the performance of the resulting 3D segmentation. Fig. 7 shows the reconstructed 3D wound segmentation of the shoulder wound, generated from 120 renderings, with color-coded true positive (light blue), false positives (blue) and false negatives (yellow). The 3D projection of 2D segmentations provides a more precise understanding of the geometric failure modes of 2D segmentation models.
@@ 5 Compliance with Ethical Standards
Fig. 7: Color-coded 3D segmentation metrics for the shoulder wound.

### v2

@@ 3.1 Baseline systems and evaluation metrics
3D wound reconstruction: A 3D reconstruction algorithm estimates the 3D geometry of an object from a collection of 2D images. The prevailing methods in the literature rely on standard projective geometry techniques such as structure-from-motion and multiview stereopsis  [23, 12, 13]. However, new deep learning approaches for 3D scene rendering (e.g. Neural Radiance Fields (NeRF) [24]), are becoming very competitive. In this paper, we conduct a comparative analysis of two prominent open-source tools for 3D reconstruction: COLMAP [25] and Meshroom [26]. We also assess the performance of NeuS-Facto, a NeRF model tailored for surface extraction from the open-source SDFStudio toolbox [24].
We compared the 3D reconstructed meshes with the ground-truth synthetic mesh, after alignment using three steps: i) align the camera positions of the ground-truth data with those estimated by the frameworks (by solving a Procrustes problem [27]); ii) crop both meshes using the ground-truth 3D mask for wound bed segmentation, followed by fine alignment using the Iterative Closest Point (ICP) algorithm (applied only to the cropped meshes); iii) apply the transformations to the original meshes, followed by cropping the wound area again to report performance on the wound area only. In Table 2, we report the Average Symmetric Distance (ASD), Hausdorff Distance (HD90), and Normal Consistency (NC) metrics.
@@
3D wound bed segmentation: We introduce a 3D wound segmentation technique that assigns binary labels to different regions of the reconstructed 3D models. We used a Meshroom-based texturing algorithm [29] to project a set of 2D wound segmentation masks onto 3D mesh vertices labeled as background and wound bed.
@@ 3.2 Results and discussion
3D wound reconstruction: Quantitative results for 3D wound reconstruction in two selected samples are reported in Table 2. In our experiment, COLMAP demonstrates superior surface accuracy, while the performance of the Neural rendering-based method is nearly comparable.
@@
3D wound segmentation: Table 4 compares 3D wound segmentation results with ground truth using previously described metrics. Notably, for the second sample, incorporating a higher number of 2D segmentation maps enhances the performance of the resulting 3D segmentation. Fig. 7 (left) shows the reconstructed 3D wound segmentation of the shoulder wound, generated from 120 renderings, with color-coded true positive (light blue), false positives (blue) and false negatives (yellow). The 3D projection of 2D segmentations provides a more precise understanding of the geometric failure modes of 2D segmentation models.
@@ 5 Compliance with Ethical Standards
Fig. 7: Color-coded 3D segmentation metrics for the shoulder wound. Right: ground-truth (red) and predicted (green) contour.